OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi/5

OpenStudy (e.mccormick):

Do you understand what coterminal is?

OpenStudy (e.mccormick):

A coterminal angle is one that is a revolution appart. For example, 0 and 2pi.

OpenStudy (e.mccormick):

So if you add 2pi, that is the positive coterminal. If you subtract, that will be negative.

OpenStudy (e.mccormick):

Remember how fractions work? You need a common denominator to add.

OpenStudy (e.mccormick):

\[2\pi\implies \frac{2\pi}{1} \cdot \frac{5}{5}\implies \frac{10\pi}{5}\]

OpenStudy (e.mccormick):

You just subtract to get the negative one. That is just what 2pi becomes to add or subtract it.

OpenStudy (e.mccormick):

Well, you have \(\frac{1\pi}{5}\) to start. I added the one for emphasis. What doi you get if you add or subtract the 10pi/5 to that?

OpenStudy (anonymous):

11 and -9pi . . . wow that's amazing! thank you so much

OpenStudy (e.mccormick):

np. Have fun! And I hope that shows you how those work. It is just a fraction. =)